OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@EmilyF03

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@campbell_st

OpenStudy (loser66):

YOu have a formula for this. Look up your notes and tell me what it is.

OpenStudy (anonymous):

i couldnt find it

OpenStudy (loser66):

Why? then google it

OpenStudy (anonymous):

\[\cos \Theta = V \times W / \left| V \right| \left| W \right|\]

OpenStudy (anonymous):

that? @Loser66

OpenStudy (loser66):

Not really. It is \(cos (\theta) =\dfrac{<v, w>}{||v||*||w||}\)

OpenStudy (loser66):

The numerator is "dot" product of v and w Hence, in your case , it is u and v. Now rewrite it, please.

OpenStudy (anonymous):

so the denominator would be 36?